# Dysca: A Dynamic and Scalable Benchmark for Evaluating Perception Ability of LVLMs

### v3

@@ 1 Introduction
Compared to previous works in Tab. 1, we provide an end-to-end process from image to Vision-QA generation. The approach significantly reduces annotation costs compared to manually labeling images (e.g., MME (Fu et al., 2023)) while achieving the correctness for evaluating LVLMs. It also avoids the risk of hallucinate annotations that may occur when using ChatGPT for labeling based on image prompts (e.g., JourneyDB (Sun et al., 2023a)). This novel pipeline enables us to create a benchmark that is easily scalable and adaptable for incorporating new subtasks and scenarios. Thanks to the generative paradigm, Dysca can be customized to meet the specific requirements of the evaluator for testing purposes.
@@ 2.1 Large Vision-Language Models
The landscape of Large Vision-Language Models (LVLMs) has been significantly shaped by the pioneering success of Large Language Models (LLMs) such as GPTs (Radford et al., 2019; Brown et al., 2020; Ouyang et al., 2022) and LLaMA (Touvron et al., 2023), catalyzing advancements in multimodal content understanding and generation (Zhang et al., 2024a), including intricate tasks like image-text comprehension. At the forefront of these developments, BLIP-2 (Li et al., 2023d) introduces a lightweight Q-Former (Li et al., 2023d) that facilitates alignment between textual and visual representations through a cross-attention mechanism (Li et al., 2023d). InstructBLIP (Dai et al., 2023) takes a step further by incorporating textual instructions into the Q-Former, which significantly improves performance. LLAVA (Liu et al., 2023b) employs GPT-4 (OpenAI, 2023) to transform data into multimodal instruction-following data and uses CLIP (Radford et al., 2021) and LLAMA (Touvron et al., 2023) for fine-tuning instructions, achieving advanced performance. LLAVA-1.5 (Liu et al., 2023a) extends this paradigm by integrating MLP projection and introducing academic task-specific Vision-language QA data. Recently, models like Otter (Li et al., 2023a), MiniGPT-4 (Zhu et al., 2023), Qwen-VL-Chat (Bai et al., 2023a) and XComposer-VL (Zhang et al., 2023) further unleash the cross-modal understanding capabilities of LVLMs. Besides, many powerful closed-source LVLMs, including Gemini-1.5-Pro (Team et al., 2024) and GPT-4o (Ope, ), have publicly released their APIs, promoting the development of downstream applications.
@@ 2.2 Benchmarks for LVLMs
The third type (Liu et al., 2023b; Fu et al., 2023; Xu et al., 2023; Shao et al., 2023; Li et al., 2023c; b; Liu et al., 2023c; Bai et al., 2023b; Li et al., 2023f; Yang et al., 2023a; Yu et al., 2023; Ge et al., 2023; Cai et al., 2023; Chen et al., 2024; Liu et al., 2024) focuses on evaluating LVLMs in an objective and large-scaled manner, where we list the detailed information of them in the Tab. 1. Some of them have been adopted by the community (Contributors, 2023) as the standard benchmarks for evaluating LVLMs (OpenAI, 2022; Zhang et al., 2023), like MME (Fu et al., 2023) and MMBench (Liu et al., 2023c). These benchmarks evaluate models through the objective answer types and most of them leverage the automatic annotation and evaluation manner for revealing the fine-grained drawbacks of current LVLMs. However, the previous benchmarks primarily concentrate on evaluating LVLMs using realistic images and clean scenario, leaving multi-stylized images and noisy scenarios unexplored. Moreover, many of them conduct QA by selecting images from publicly available datasets (e.g., (Lin et al., 2014; Russakovsky et al., 2014)). While they state that the questions have been re-annotated, they cannot guarantee that the LVLMs have not seen the image during training stage. The previous work (Chen et al., 2024) has proved that these benchmarks have unintentionally leaked into the training data of LLMs and LVLMs. One possible way to solve data leakage is using novel but synthesis images, where JourneyDB (Sun et al., 2023a) is the first work aiming to leverage synthesis images to evaluate current LVLMs. The prompts and the corresponding images are downloaded from Midjourney (mid, ) and ChatGPT (OpenAI, 2022) is leveraged to label the images. However, JourneyDB is a top-down framework where the number of images is fixed. Besides, the ChatGPT labeling may cause hallucinate annotations, leading to the unreliable evaluation results. In contrast, our Dysca serves as the bottom-up framework, allowing for dynamic and scalable generation for both images and evaluation questions and involving more evaluation tasks. The rule-based question generation method also makes the annotations more accuracy.
@@ 3.2 Perceptual Tasks
Evaluation dimensions. Perception is one of the most fundamental capabilities of LVLMs and previous works (Fu et al., 2023) have shown that the lack of perceptual ability may result in hallucination (Li et al., 2023e). In order to comprehensively evaluate LVLMs’ perception capability, we collect and organize existing sub-dimensions from current evaluation datasets, resulting in a total of 20 assessment dimensions where we show all the subtasks and the corresponding amount of their annotation in the Fig. 2. We investigate on two types of perception dimensions, i.e., coarse-grained and fine-grained perception. Coarse-grained perception involves recognizing the style, background and color of images. Fine-grained perception involves recognizing the animal, object, plant, food, age, gender, expression, race, celebrity, action, text, clothes, movie, anime, landmark, profession and TV shows.
@@ 3.3 Construction of Questions & Answers
Recall that the data generation for Dysca benchmark consists of four dimensions, i.e., $(M,P,I,Q)$, denoting the metadata ($M$), prompt ($P$), image ($I$) and question-answer pairs ($Q$), respectively. The relationships between these parts and the process of constructing Dysca are shown in Fig. 3. The metadata M is the core of the whole Dysca, containing all the information for generating $P$, $I$ and $Q$. The metadata $M$ consists of foreground, attribute, background and style, and these information guide the generation of the prompt ($P$) through pre-designed templates. Then, we utilize the T2I diffusion model to generate the corresponding image using the prompt $P$. For generating the image with specific text on it for the OCR subtask, we leverage TextDiffusion2 (Chen et al., 2023a), which is the state-of-the-art text rendering method. For the rest of images, we leverage Stable Diffusion XL (Podell et al., 2023). Subsequently, based on the different question types we select, i.e., multi-choices, true-or-false and free-form, we generate the corresponding VQA pairs in Dysca.
@@
Figure 3: The process of generating the prompt (P), image (I) and question-answer pairs (Q) from the metadata (M).
@@ 3.4 Evaluation Strategy
Besides, thanks to generative evaluation framework of Dysca, we are able to effectively control variables and conduct a detailed analysis of the model’s fine-grained capabilities. Specifically, we introduce two novel metrics to measure the sensitivity of LVLMs on question types and covariate shift. The sensitivity to question type aims to evaluate whether LVLMs exhibit inconsistent performance when facing different types of question types (i.e., multi-choice vs. true-or-false). We first normalize the score of true-or-false by $TFSQ=\frac{S-50}{100-50}*100\%$ and multi-choices by $MCSQ=\frac{S-25}{100-25}*100\%$, where $S$ denotes the score of the models in each question type. The sensitivity of LVLMs to question types is defined by:
@@
The sensitivity to covariate shift aims to evaluate whether LVLMs exhibit inconsistent performance when facing the same content and question format, but with variations in image covariates (i.e., image style). It is defined by:
@@
where $SC_{Avg}=\frac{\sum_{i=1}^{N}S}{N}$ and $S_{i}$ denotes the score of the models in each style. Since there are 51 image styles in Dysca, we set $N=51$.
@@ 4.2 Key Observations
(1) For LVLMs, the capacity of the language model plays a crucial role. When using the same visual encoder, models that adopt a language model with larger parameter sizes (e.g., Vicuna-13B models generally outperform Vicuna-7B models by 8%) or models with a different but stronger architecture (e.g., GLM-4V (GLM et al., 2024) with GLM-4-9B-Chat shows a 20% performance increase compared to Emu2 (Sun et al., 2024) with LLaMA-33B) tend to achieve better performance.
@@
(2) Each model shows robustness in the corruption scenario, but experiences significant degradation in both attack scenarios. In the image corruption scenario, all models demonstrate minimal score variations (less than 1%), indicating consistent performance under non-targeted disruptions. However, under print attacks, performance drops are notable. For example, two closed-source models experience substantial declines: Gemini-1.5-pro (Team et al., 2024) drops by 21.5% (from 77.79 to 61.05), and GPT-4o (OpenAI, 2023) suffers a 25.8% decrease (from 75.69 to 56.10). Among the advanced open-source models, GLM-4V’s (GLM et al., 2024) performance decreases sharply by 36.5% (from 82.09 to 52.09), and Phi-3-Vision (Abdin et al., 2024) records a 21.39% drop (from 73.23 to 57.78). Notably, mPLUG-Owl-2 (Ye et al., 2023) shows the highest robustness, with only a 1.7% reduction (from 74.09 to 72.84), while the XComposer-VL (Zhang et al., 2023; Dong et al., 2024) series also exhibits strong resistance to print attacks. In the adversarial attack scenario, where the attack algorithm directly targets the image encoder, LVLMs with shared encoder architecture (e.g., Blip2 (Li et al., 2023d), InstructBLIP (Dai et al., 2023), and XComposer-VL (Zhang et al., 2023), all using EVA-CLIP (Fang et al., 2022) as their image encoder) suffer significant performance drops, with some models even performing worse than random chance. For instance, XComposer-VL (Zhang et al., 2023) experiences a 57.6% drop (from 71.40 to 30.28). Models with alternative image encoders also experience degradation, ranging from 1% to 5%. The greater impact of adversarial noise compared to corruption noise suggests that adversarial attacks exhibit a certain degree of transferability across different model architectures. More detailed results are available in Appendix F.
Figure 5: The Visualization of hierarchical clustering results of 20 subtasks.
Figure 6: The Visualization of hierarchical clustering results of 51 animal categories.
@@
(3) Models exhibit varying sensitivity to different question types and covariate shifts. For the Sensitivity to Question Types, Xcomposer2-VL (Dong et al., 2024) achieves the best result with a score of 20.9. However, we observe that the perception ability of LVLM does not show a positive correlation with sensitivity to question types. For instance, while the GLM-4V (GLM et al., 2024) achieves the highest performance in evaluation tasks, it exhibits higher sensitivity to question types than Xcomposer2-VL (Dong et al., 2024). One of the factors influencing the sensitivity to question types may be the inherent biases of the language model. Using the same base language model may result in similar outcomes for this metric. it is also noted that Gemini (Team et al., 2024) performs bad in this metric, revealing its preference for certain question types. For the Sensitivity to Covariate Shifts, GLM-4V (Dong et al., 2024) achieves the best result with a score of 0.8. However, we also observe that the perception ability of LVLM does not show a positive correlation with covariate shifts. For example, InstructBLIP-Flan-T5-XXL (Dai et al., 2023) outperforms InstructBLIP-Flan-T5-XL in terms of performance but shows higher sensitivity to covariate shifts.
@@ 4.3 Analysis on inter-task and intra-task
Analysis on Inter-task. Here, we perform hierarchical clustering based on the Euclidean distance of the scores across 20 evaluation dimensions for 26 models. As shown in the Fig. 6, based on the consistency of model performance across the 20 evaluation dimensions, the models can be categorized into four tiers. We also observe that LVLMs tend to perform better in dimensions that involve well-defined image perception, such as landmark recognition and object recognition. However, dimensions likes style recognition and movie recognition exhibit greater challenges for LVLMs, which may be attributed to the limited training resources in these specific domains. Besides, commercial models exhibit poor performance in tasks related to “people” (i.e., “race”, “age”, and “gender”). This is likely due to the additional safety training incorporated into closed-source models, which results in more conservative responses on related questions. Detailed results are provided in Appendix A.
Analysis on Intra-task. Thanks to the diverse metadata design, Dysca enables highly granular analysis within a single task. Taking animal categories as an example, we analyze the performance of LVLMs on different animals. Here, we perform same hierarchical clustering methods across 51 animal categories. As shown in the Fig. 6, LVLMs performance vary significantly across these categories. Models tend to perform poorly when applied to marine life. This could be caused by the challenges in collecting ocean-related data or possibly to the fact that human understanding of land environments is better than that of the ocean. This observation highlights the need to direct the model’s focus to oceanic domains.
@@ 4.4 The Validity of Dysca
In this section, we investigate on the evaluation gap between Dysca and non-synthesis benchmarks. We calculate the Spearman’s rank correlation coefficient (Spearman, 1904) $\rho$ and the Kendall rank correlation coefficient (KENDALL, 1938) $\tau$ between the evaluation ranking of Dysca under clean scenario with the non-synthesis benchmark’s evaluation ranking, i.e., MMBench (Liu et al., 2023c), OCRBench (Liu et al., 2024) and SeedBench-2 (Li et al., 2023b). Both coefficient generate a score in the range of [-1,1], where 1 represents a perfect positive correlation, -1 represents a perfect negative correlation, and 0 represents no correlation. These coefficients are typical tools for measuring the correlation between variables in statistics. When the absolute value of either coefficient exceeds 0.6, it is considered to indicate a significant correlation  (Akoglu, 2018). Specifically, we intersect our Dysca with current benchmarks based on the perceptual subtasks, evaluation models and evaluation question types. We then calculate the correlation of model evaluation rankings within this intersection. The results are shown in the first row of Tab. 4. For the MMbench (Liu et al., 2023c) and OCRBench (Liu et al., 2024), both metrics show the high correlation, with $\rho$ and $\tau$ higher than 0.6. However, the correlation for SeedBench-2 (Li et al., 2023b) is not as strong. Considering that SeedBench-2 only contains realistic images, we conduct additional experiments using the evaluation ranks on our realistic style images only. As shown in the second row of Tab. 4, the correlation results of SeedBench-2 significantly improve (i.e., 0.46 vs. 0.64 for $\rho$ and 0.43 vs. 0.62 for $\tau$). The correlation with OCRBench also improves to 1, demonstrating the validity of using synthetic datasets for evaluation LVLMs.
@@
Besides, we calculate the data distribution distance between each benchmark to prove the low distance distribution between Dysca and non-synthesis benchmarks. We select CCBench (Liu et al., 2023c), COCO-Val  (Lin et al., 2014), MMVet (Yu et al., 2023), MMBench (Liu et al., 2023c), MME (Fu et al., 2023), MMStar (Chen et al., 2024), OCRBench (Liu et al., 2024) and ScienceQA (Lu et al., 2022). The reason why we choose these benchmark is that they have been widely used in evaluating LVLMs. We use Kernel Maximum Mean Discrepancy (KMMD) (Schölkopf et al., 2007) to measure the distribution distance. Specifically, we randomly sample 3,000 images from each benchmark (if the scale of the benchmark less than 3000, we use all that data) and utilize CLIP  (Radford et al., 2021) to encode these images. Then, we calculate the KMMD value using an RBF kernel between each pair. The results are shown in Fig. 7, with darker colors indicating larger distances. Each row represents the value of KMMD between two benchmarks. The value of last row denotes the average value of KMMD. As can be seen, the distribution distance between Dysca and real-image benchmarks ranks in the middle compared to all other benchmarks. This demonstrates that the distribution between our benchmark and real benchmarks is similar, indicating that the evaluation results can effectively reflect the model’s performance in real-world scenarios.

### v4

@@ 1 Introduction
Compared to previous works in Tab. 1, we provide an end-to-end process from image to Vision-QA generation. This approach significantly reduces annotation costs compared to manually labeling images (e.g., MME (Fu et al., 2023)) while achieving the correctness for evaluating LVLMs. It also avoids the risk of hallucinate annotations that may occur when using ChatGPT for labeling based on image prompts (e.g., JourneyDB (Sun et al., 2023a)). This novel pipeline enables us to create a benchmark that is easily scalable and adaptable for incorporating new subtasks and scenarios. Thanks to the generative paradigm, Dysca can be customized to meet the specific requirements of the evaluator for testing purposes.
@@ 2.1 Large Vision-Language Models
The landscape of Large Vision-Language Models (LVLMs) has been significantly shaped by the pioneering success of Large Language Models (LLMs) such as GPTs (Radford et al., 2019; Brown et al., 2020; Ouyang et al., 2022) and LLaMA (Touvron et al., 2023), catalyzing advancements in multimodal content understanding and generation (Zhang et al., 2024a), including intricate tasks like image-text comprehension. At the forefront of these developments, BLIP-2 (Li et al., 2023d) introduces a lightweight Q-Former (Li et al., 2023d) that facilitates alignment between textual and visual representations through a cross-attention mechanism (Li et al., 2023d). InstructBLIP (Dai et al., 2023) takes a step further by incorporating textual instructions into the Q-Former, which significantly improves the zero-shot performance. LLAVA (Liu et al., 2023b) employs GPT-4 (OpenAI, 2023) to transform data into multimodal instruction-following data and uses CLIP (Radford et al., 2021) and LLAMA (Touvron et al., 2023) for fine-tuning instructions, achieving advanced multimodal chat abilities. LLAVA-1.5 (Liu et al., 2023a) extends this paradigm by integrating MLP projection and introducing academic task-specific Vision-language QA data. Recently, models like Otter (Li et al., 2023a), MiniGPT-4 (Zhu et al., 2023), Qwen-VL-Chat (Bai et al., 2023a) and XComposer-VL (Zhang et al., 2023) further unleash the cross-modal understanding capabilities of LVLMs. Besides, many powerful closed-source LVLMs, including Gemini-1.5-Pro (Team et al., 2024) and GPT-4o (Ope, ), have publicly released their APIs, promoting the development of downstream applications.
@@ 2.2 Benchmarks for LVLMs
The third type (Liu et al., 2023b; Fu et al., 2023; Xu et al., 2023; Shao et al., 2023; Li et al., 2023c; b; Liu et al., 2023c; Bai et al., 2023b; Li et al., 2023f; Yang et al., 2023a; Yu et al., 2023; Ge et al., 2023; Cai et al., 2023; Chen et al., 2024; Liu et al., 2024) focuses on evaluating LVLMs in an objective and large-scaled manner, where we list the detailed information of them in the Tab. 1. Some of them have been adopted by the community (Contributors, 2023) as the standard benchmarks for evaluating LVLMs (OpenAI, 2022; Zhang et al., 2023), like MME (Fu et al., 2023) and MMBench (Liu et al., 2023c). These benchmarks evaluate models through the objective answer types and most of them leverage the automatic annotation and evaluation manner for revealing the fine-grained drawbacks of current LVLMs. However, the previous benchmarks primarily concentrate on evaluating LVLMs using realistic images under clean scenario, leaving multi-stylized images and noisy scenarios unexplored. Moreover, many of them conduct QA by selecting images from publicly available datasets (e.g., (Lin et al., 2014; Russakovsky et al., 2014)). While they state that the questions have been re-annotated, they cannot guarantee that the LVLMs have not seen the image during training stage. The previous work (Chen et al., 2024) has proved that these benchmarks have unintentionally leaked into the training data of LLMs and LVLMs. One possible way to solve data leakage is using novel but synthesis images, where JourneyDB (Sun et al., 2023a) is the first work aiming to leverage synthesis images to evaluate current LVLMs. The prompts together with the corresponding images are downloaded from Midjourney (mid, ) and ChatGPT (OpenAI, 2022) is leveraged to label the images. However, JourneyDB is a top-down framework where the number of images is fixed. Besides, the ChatGPT labeling may cause hallucinate annotations, leading to the unreliable evaluation results. In contrast, Dysca serves as a bottom-up framework which enables dynamic and scalable generation of both images and evaluation questions, while also supporting varied evaluation tasks. The rule-based question generation method also makes the annotations more accuracy.
@@ 3.2 Perceptual Tasks
Evaluation dimensions. Perception is one of the most fundamental capabilities of LVLMs and previous works (Fu et al., 2023) have shown that the lack of perceptual ability may result in hallucination (Li et al., 2023e). In order to comprehensively evaluate LVLMs’ perception capability, we collect and organize existing sub-dimensions from current benchmarks, resulting in 20 assessment dimensions where we show all the subtasks and the corresponding amount of their annotation in the Fig. 2. We investigate on two types of perception dimensions, i.e., coarse-grained and fine-grained perception. Coarse-grained perception involves recognizing the style, background and color of images. Fine-grained perception involves recognizing the animal, object, plant, food, age, gender, expression, race, celebrity, action, text, clothes, movie, anime, landmark, profession and TV shows.
@@ 3.3 Construction of Questions & Answers
Recall that the data generation for Dysca benchmark consists of four dimensions, i.e., $(M,P,I,Q)$, denoting the metadata ($M$), prompt ($P$), image ($I$) and question-answer pairs ($Q$), respectively. The relationships between these parts and the process of constructing Dysca are shown in Fig. 3. The metadata M is the core of the whole Dysca, containing all the information for generating $P$, $I$ and $Q$. The metadata $M$ consists of foreground, attribute, background and style, which guides the generation of the prompt ($P$) through pre-designed templates. Then, we utilize the T2I diffusion model to generate the corresponding image using the prompt $P$. For generating the image with a specific text on it for the OCR subtask, we leverage TextDiffusion2 (Chen et al., 2023a), which is the state-of-the-art text rendering method. For the rest of images, we leverage Stable Diffusion XL (Podell et al., 2023). Subsequently, based on the different question types we select, i.e., multi-choices, true-or-false and free-form, we generate the corresponding VQA pairs in Dysca.
@@
Figure 3: The process of generating the prompt (P), image (I) and QA pairs (Q) from metadata (M).
@@ 3.4 Evaluation Strategy
Figure 5: Models exhibit different performance when facing the same image but different question types.
@@
Besides, thanks to generative evaluation framework of Dysca, we are able to effectively control variables and conduct a detailed analysis of the model’s fine-grained capabilities. Specifically, we introduce two novel metrics to measure the sensitivity of LVLMs on question types and covariate shift (i.e., image style). The sensitivity to question type aims to evaluate whether LVLMs exhibit inconsistent performance when facing different types of question types (i.e., multi-choice vs. true-or-false). We first normalize the score of true-or-false by $TFSQ=\frac{S-50}{100-50}*100\%$ and multi-choices by $MCSQ=\frac{S-25}{100-25}*100\%$, where $S$ denotes the score of the models in each question type. The sensitivity of LVLMs to question types is defined by:
@@
The sensitivity to covariate shift aims to evaluate whether LVLMs exhibit inconsistent performance when facing the same content and question format, but with variations in image covariates. It is defined by:
@@
where $SC_{Avg}=\frac{\sum_{i=1}^{N}S_{i}}{N}$ and $S_{i}$ denotes the score of the models in each style. Since there are 51 image styles in Dysca, we set $N=51$.
@@ 4.2 Key Observations
(1) For LVLMs, the capacity of the language model plays a crucial role. When using the same visual encoder, models that utilize a language model with a larger parameter sizes (e.g., MiniGPT-4 achieves an improvement score of 8.79 in the clean scenario when using Vicuna-13B compared to Vicuna-7B) or with a stronger capability (e.g., GLM-4V that uses the GLM-4-9B-Chat language model shows an improvement score of 18.45 in the clean scenario compared to Emu2, which uses the LLaMA-33B language model) tend to achieve a better performance.
@@
(2) Each model shows robustness in the corruption scenario, but experiences significant degradation in both attack scenarios. In the image corruption scenario, all models demonstrate minimal score variations, i.e., less than 1%. However, under print attacks, performance drops are notable. For instance, two closed-source models exhibit significant performance degradation: Gemini-1.5-pro declines by 21.5%, resulting in a score reduction from 77.79 to 61.05, while GPT-4o shows a 25.8% decrease, dropping its score from 75.69 to 56.10. Among the leading open-source models, GLM-4V experiences a sharp 36.5% performance drop, lowering its score from 82.09 to 52.09, and Phi-3-Vision records a 21.39% decline, reducing its score from 73.23 to 57.78. Notably, mPLUG-Owl-2 demonstrates the highest robustness, with only a 1.7% reduction. The XComposer-VL series also exhibits strong resilience against print-based attacks. In the adversarial attack scenario, where the attack algorithm directly targets the image encoder, LVLMs employing a shared encoder architecture (e.g., Blip2, InstructBLIP, and XComposer-VL, all of which utilize EVA-CLIP (Fang et al., 2022) as their image encoder) suffer substantial performance declines, with some models performing even worse than random chance. For example, XComposer-VL experiences a 57.6% drop, reducing its score from 71.40 to 30.28. Models with different image encoders also experience performance degradation ranging from 1% to 5%, showing a higher impact than that of corruption noise. More detailed results are available in Appendix F.
@@
(3) Models exhibit varying sensitivity to different question types and covariate shifts. For the sensitivity to question types, we present two examples in Fig. 5. As can be seen, the XComposer-VL (Zhang et al., 2023) recognizes the sparrow in the image under a multiple-choice setting but fails to identify the sparrow in the same image under a true-or-false setting. The quantity results are shown in Tab. 3. Xcomposer2-VL achieves the best result with a score of 20.9. However, we observe that the perception ability of LVLM does not show a positive correlation with sensitivity to question types. For instance, while the GLM-4V achieves the highest performance in evaluation tasks, it exhibits higher sensitivity to question types than Xcomposer2-VL. One of the factors influencing the sensitivity to question types may be the inherent biases of the language model. Using the same base language model may result in similar outcomes for this metric. it is also noted that Gemini-1.5-pro performs bad in this metric, revealing its preference for certain question types. For the sensitivity to covariate shifts, as shown in the last column of Tab. 3, GLM-4V achieves the best result with a score of 0.8. However, we also observe that the perception ability of LVLM does not show a positive correlation with covariate shifts. For example, InstructBLIP-Flan-T5-XXL outperforms InstructBLIP-Flan-T5-XL in terms of performance but shows higher sensitivity to covariate shifts.
@@ 4.3 Analysis on inter-task and intra-task
Analysis on Inter-task. In order to investigate the inter-relationships across evaluation dimensions, we conduct hierarchical clustering based on the Euclidean distance of the scores across 20 dimensions for 26 models. We observe that models show varied consistency of performance across the dimensions. LVLMs tend to perform better in dimensions that involve well-defined image perception, such as landmark recognition and object recognition. However, dimensions likes style recognition and movie recognition exhibit greater challenges for LVLMs, which may be attributed to the limited training resources in these specific domains. Besides, commercial models exhibit poor performance in tasks related to “people” (i.e., “race”, “age”, and “gender”). This is likely due to the additional safety training incorporated into closed-source models, which results in more conservative responses on related questions. Detailed results are provided in Appendix A.
Analysis on Intra-task. Thanks to the diverse metadata design, Dysca enables highly granular analysis within a single task. Taking animal categories as an example, we analyze the performance of LVLMs on different animals. Here, we perform the same hierarchical clustering methods across 51 animal categories. LVLMs performance vary significantly across these categories. We find that models tend to perform poorly when applied to marine life. This could be caused by the challenges in collecting ocean-related data. This observation highlights the need to direct the model’s focus to oceanic domains.
@@ 4.4 The Validity of Dysca
In this section, we investigate on the evaluation gap between Dysca and non-synthesis benchmarks. We calculate the Spearman’s rank correlation coefficient (Spearman, 1904) $\rho$ and the Kendall rank correlation coefficient (KENDALL, 1938) $\tau$ between the evaluation ranking of Dysca under clean scenario with the non-synthesis benchmark’s evaluation ranking, i.e., MMBench (Liu et al., 2023c), OCRBench (Liu et al., 2024) and SeedBench-2 (Li et al., 2023b). Both coefficient generate a score in the range of [-1,1], where 1 represents a perfect positive correlation, -1 represents a perfect negative correlation, and 0 represents no correlation. These coefficients are typical tools for measuring the correlation between variables in statistics. When the absolute value of either coefficient exceeds 0.6, it is considered to indicate a significant correlation  (Akoglu, 2018). Specifically, we intersect our Dysca with current benchmarks based on the perceptual subtasks, evaluation models and evaluation question types. We then calculate the correlation of model evaluation rankings within this intersection. The results are shown in the first row of Tab. 4. For the MMbench and OCRBench, both metrics show the high correlation, with $\rho$ and $\tau$ higher than 0.6. However, the correlation for SeedBench-2 is not as strong. Considering that SeedBench-2 only contains realistic images, we conduct additional experiments using the evaluation ranks on our realistic style images only. As shown in the second row of Tab. 4, the correlation results of SeedBench-2 significantly improve (i.e., 0.46 vs. 0.64 for $\rho$ and 0.43 vs. 0.62 for $\tau$). The correlation with OCRBench also improves to 1, demonstrating the validity of using synthetic datasets for evaluation LVLMs.
@@
Besides, we calculate the data distribution distance between each benchmark to prove the low distance distribution between Dysca and non-synthesis benchmarks. We select CCBench (Liu et al., 2023c), COCO-Val  (Lin et al., 2014), MMVet (Yu et al., 2023), MMBench (Liu et al., 2023c), MME (Fu et al., 2023), MMStar (Chen et al., 2024), OCRBench (Liu et al., 2024) and ScienceQA (Lu et al., 2022). The reason why we choose these benchmark is that they have been widely used in evaluating LVLMs. We use Kernel Maximum Mean Discrepancy (KMMD) (Schölkopf et al., 2007) to measure the distribution distance. Specifically, we randomly sample 3,000 images from each benchmark (if the scale of the benchmark less than 3000, we use all that data) and utilize CLIP  (Radford et al., 2021) to encode these images. Then, we calculate the KMMD value using an RBF kernel between each pair. The results are shown in Fig. 6. Each row represents the value of KMMD between two benchmarks. The value of last row denotes the average value of KMMD. As can be seen, the distribution distance between Dysca and real-image benchmarks ranks in the middle compared to all other benchmarks, indicating that the evaluation results can effectively reflect the model’s performance in real-world scenarios.